OpenStudy (anonymous):

The slope of a line is 1/8 . The point (8, 9) is on the line. Determine the equation of this line.

OpenStudy (amistre64):

there are many equation to define this line; the most common tho is the slope intercept... y = 1/8 (x-8) +9 y = (1/8)x -1+9 y = (1/8)x +8

OpenStudy (anonymous):

u should answer all my questions lol